# Consistency of Neural Causal Partial Identification

### v1

@@ 3.2 Approximating Mixed Distributions by Deep Neural Networks
Assumption 5 (Lower Bound).
@@
5 implies that $d\lambda/d(f^{-1}_{\#}\mathbb{P})$ exists and is lowered bounded by a constant $C_{f}$. The next proposition extends the Skorohod representation theorem [47]. It shows that under 5, it is possible to simulate any distribution on the unit cubes with Hölder continuous curves and uniform distribution on $[0,1]$.
Proposition 1.
Let $\lambda$ be the Lebesgue measure on $[0,1]$. Given any probability measure $\mathbb{P}$ that satisfies 5, there exists a continuous curve $\gamma:[0,1]\rightarrow\text{supp}(\mathbb{P})$ such that $\gamma_{\#}\lambda=\mathbb{P}$. Furthermore, if $d\geqslant 1$, $\gamma$ is $1/d$-Hölder continuous.
Results from [52] show that we can approximate any Hölder continuous $d$-dimensional function with index $\alpha$ by a deep ReLu network with depth $L$ and error $O(L^{-2\alpha/d})$. Leveraging this result, we can replace the first part of the architecture in the previous subsection with deep ReLu networks (See Figure 5 in the appendix). The remaining two parts are the same as the construction in Figure 1(a). The following theorem gives a sharper bound on the approximation error compared with Theorem 2.
Theorem 3.
Under the 4, if in addition $\mathbb{P}$ constrained to each component satisfies 5, there exists a neural network $\hat{g}=\hat{g}_{3}^{\tau}\circ\hat{g}_{2}\circ{\hat{g}_{1}}$ with the above architecture such that
$W(\hat{g}_{\#}\lambda,\mathbb{P})\leqslant O\left(L_{1}^{-2/\max_{i}\{d_{i}^{C}\}}+L_{2}^{-2/\max_{i}\{d_{i}^{C}\}}+(\tau-\tau\log\tau)\right),$
where $d_{i}^{C}$ are the dimensions of the connected components in 4, $\hat{g}_{i},i=1,2$ has the form $\left(\hat{g}_{i}^{1}(x_{1}),\cdots,\hat{g}_{i}^{N_{C}}(x_{N_{C}})\right)$ and $\hat{g}_{1}^{j}\in\mathcal{NN}_{1,d_{j}^{C}}(\Theta(d_{j}^{C}),L_{1})$. $\hat{g}_{2}^{j},\hat{g}_{3}^{\tau},\tau$ are the same as in Theorem 2.
Let $N$ be the number of nonzero weights in a neural network, Theorem 3 shows that a DNN can achieve $O(N^{-2/d})$ error while by Theorem 2 a wide network can only achieve $O(N^{-1/d})$ error. Now, we can put things together to construct NCM approximations. For simplicity, we omit the input and output dimensions of the neural network. As we mention in previous sections, our construction (5) consists of two parts, $\hat{f}_{i}$ approximating the structural functions $f_{i}$ in (3), and $\hat{g}_{j}(Z_{j})=\hat{g}_{3,j}^{\tau}\circ\hat{g}_{2,j}\circ\hat{g}_{1,j}(Z_{j})$ approximating the latent variables $U_{j}$. Let $\text{NCM}_{\mathcal{G}}(\mathcal{F}_{0},\mathcal{F}_{1},\mathcal{F}_{2},\tau)$ be a collection of NCMs with structural equation (5) and
$\hat{f}_{i}\in\mathcal{F}_{0},\hat{g}_{1,j}=(\hat{g}^{1}_{1,j},\cdots,\hat{g}^{N_{C,j}}_{1,j}),\hat{g}^{i}_{1,j}\in\mathcal{F}_{1},\hat{g}_{2,j}=(\hat{g}^{1}_{2,j},\cdots,\hat{g}^{N_{C,j}}_{2,j}),\hat{g}^{i}_{2,j}\in\mathcal{F}_{2},$
@@
Corollary 1.
Given a causal model $\mathcal{M}^{*}\in\mathcal{M}(\mathcal{G},\mathcal{F}_{L},\boldsymbol{U})$, suppose that Assumptions 1-3 hold and the distributions of $U_{C}$ for all $C^{2}$-component satisfy the assumptions of Theorem 3. Let $d_{\max}^{\text{in}}$ and $d_{\max}^{\text{out}}$ be the largest input and output dimension of $f_{i}$ in (3) and $d_{\max}^{U}$ be the largest dimension of all latent variables. There exists a neural causal model
^$\hat{\mathcal{M}}\in\text{NCM}_{\mathcal{G}}(\mathcal{NN}(\Theta(d^{\text{in}}_{\max}d^{\text{out}}_{\max}),L_{0}),\mathcal{NN}\left(\Theta\left(d^{U}_{\max}\right),L_{1}\right),\mathcal{NN}\left(\Theta\left((d^{U}_{\max})^{2}\right),L_{2}\right),\tau)$
with structure equations (5). For any intervention $\boldsymbol{T}=\boldsymbol{t}$, $\hat{\mathcal{M}}$ satisfies
$W\left(P^{\mathcal{M}^{*}}(\boldsymbol{V}(\boldsymbol{t})),P^{\hat{\mathcal{M}}}(\boldsymbol{V}(\boldsymbol{t}))\right)\leqslant O(L_{0}^{-2/d^{\text{in}}_{\max}}+L_{1}^{-2/d^{U}_{\max}}+L_{2}^{-2/d^{U}_{\max}}+(\tau-\tau\log\tau)).$
@@ 4 Consistency of Neural Causal Partial Identification
In this section, we prove the consistency of the max/min optimization approach to partial identification. In the finite sample setting, we consider the following estimator of the optimal values of (1).
$\min_{\hat{\mathcal{M}}\in\text{NCM}_{\mathcal{G}}(\mathcal{F}_{0,n},\mathcal{F}_{1,n},\mathcal{F}_{2,n},\tau_{n})}$ $\mathbb{E}_{t\sim\mu_{T}}\mathbb{E}_{\hat{\mathcal{M}}}[F(V_{1}(t),\cdots,V_{n_{V}}(t))],$ (6)
$s.t.~{}~{}~{}$ $S_{\lambda_{n}}(P_{m_{n}}^{\hat{\mathcal{M}}}(\boldsymbol{V}),P^{\mathcal{M}^{*}}_{n}(\boldsymbol{V}))\leqslant\alpha_{n},$
where $P^{\mathcal{M}^{*}}_{n},P_{m_{n}}^{\hat{\mathcal{M}}}$ are the empirical distribution of $P^{\mathcal{M}^{*}},P^{\hat{\mathcal{M}}}$ with sample size $n,m_{n}$, $\mu_{T}$ is some reference measure and $\mathcal{F}_{i,n}$ will be specified later. For example, the ATE is a special case of the objective. Let $F_{n}$ be the solution of (6). We use the Sinkhorn distance here because it can be computed efficiently in practice [15]. We want to study if $F_{n}$ gives a useful lower bound as $n\rightarrow\infty$.
@@
Proposition 2 (Informal, see Proposition 5 for a formal version).
There exists a constant $c>0$ and an identifiable SCM $\mathcal{M}^{*}$ satisfying Assumptions 1-5 such that for any $\epsilon>0$, there exists an SCM $\mathcal{M}_{\epsilon}$ satisfying $W(P^{\mathcal{M}^{*}}(\boldsymbol{V}),P^{\mathcal{M}_{\epsilon}}(\boldsymbol{V}))\leq\epsilon$ and $|\text{ATE}_{\mathcal{M}^{*}}-\text{ATE}_{\mathcal{M}_{\epsilon}}|>c$.
Here, $\mathcal{M}^{*}$ is the ground-truth model and $\mathcal{M}_{\epsilon}$ are the models we use to approximate $\mathcal{M}^{*}$. Proposition 2 implies that we need some regularization on $\mathcal{M}_{\epsilon}$. Otherwise, even if the observation distributions are close, their ATEs can be far away. In particular, we may want to regularize the Lipschitz constant of the NN. Much work has been done to impose Lipschitz regularization during the training process [13, 49, 22, 41, 10]. We denote $\text{Lip}(f)$ to be the Lipschitz constant of a function $f$ and define the truncated Lipschitz NN class,
@@
For simplicity, we omit the dimensions and use shorthand $\mathcal{N}\mathcal{N}^{L_{f},K}(W,L)$. The next theorem gives the consistency result of the min estimator. To state the theorem, we define $F_{*}=\mathbb{E}_{t\sim\mu_{T}}\mathbb{E}_{\mathcal{M}^{*}}[F(V_{1}(t),\cdots,V_{n_{V}}(t))]$ to be the true value and $\underline{F}^{L}$ to be the optimal value of the optimization problem over SCMs with Lipschitz constant $L$.
$\min_{\hat{\mathcal{M}}\in\mathcal{M}\left(\mathcal{G},\mathcal{F}^{K}_{L},\boldsymbol{U}\right),P(\boldsymbol{U})}$ $\mathbb{E}_{t\sim\mu_{T}}\mathbb{E}_{\hat{\mathcal{M}}}[F(V_{1}(t),\cdots,V_{n_{V}}(t))],$ (7)
$s.t.$ $W(P^{\hat{\mathcal{M}}}(\boldsymbol{V}),P^{\mathcal{M}^{*}}(\boldsymbol{V}))=0,$
where the minimum is taken over $\mathcal{M}\left(\mathcal{G},\mathcal{F}^{K}_{L},\boldsymbol{U}\right)$ with $\mathcal{F}^{K}_{L}=\{f:\|f\|\leqslant K,f\in\mathcal{F}_{L}\}$ and all latent distributions $P(\boldsymbol{U})$. Note that if $L$ is the Lipschitz bound $L_{f}$ that we assume on our structural functions, then $\underline{{F}}^{L}$ is equal to the sharp lower bound $\underline{{F}}$.
Theorem 4.
Let $\mathcal{M^{*}}$ be any SCM satisfying the assumptions of Corollary 1. Suppose that the Lipschitz constant of functions in $\mathcal{M}$ is $L_{f}$, $F:\mathbb{R}^{n_{V}}\rightarrow\mathbb{R}$ in (6) is Lipschitz continuous and $\tau_{n}>0$, let $K>0$ be the constant in 3, $\hat{L}_{f}=\sqrt{d_{\max}^{\text{in}}d_{\max}^{\text{out}}}L_{f}$,
$\mathcal{F}_{0,n}=$ $\mathcal{NN}^{\hat{L}_{f},K}\left(W_{0,n},\Theta\left(\log d^{\text{in}}_{\max}\right)\right),\mathcal{F}_{1,n}=\mathcal{NN}^{\infty,\infty}(\Theta(d^{U}_{\max}),L_{i,n}),$
@@
where $d_{\max}^{U},d_{\max}^{\text{in}}$ and $d_{\max}^{\text{out}}$ are defined in Corollary 1, take the radius to be $\alpha_{n}=\epsilon_{n}+s_{n},\epsilon_{n}=O(W_{0,n}^{-1/d^{\text{in}}_{\max}}+\sum_{i=1}^{2}L_{i,n}^{-2/d^{U}_{\max}}+\tau_{n}\log\tau_{n}),s_{n}=O(m_{n}^{-1/\left(d_{\max}^{U}+2\right)}\log m_{n}+\delta_{n}+\log(nm_{n})\lambda_{n})$ and $\delta_{n}=O(n^{-1/{\max\{2,d_{\max}^{U}\}}}\log^{2}(n))$. If $m_{n}=\Omega(n),L_{i,n}=\Theta(m_{n}^{d_{\max}^{U}/\left(2d_{\max}^{U}+4\right)}),i=1,2$ and $\lim_{n\rightarrow\infty}\min\{\tau_{n}^{-1},W_{0,n},(\log(nm_{n})\lambda_{n})^{-1}\}=\infty$, then with probability 1, $[\liminf_{n\rightarrow\infty}F_{n},\limsup_{n\rightarrow\infty}F_{n}]\subset[\underline{F}^{\hat{L}_{f}},F_{*}]$.
Note that our theorem shows that the solution converges to the point $\underline{{F}}^{\hat{L}_{f}}$, where $\hat{L}_{f}$ is slightly larger than he original constraint $L_{f}$ we impose on the structural functions. Hence, this point can potentially be slightly smaller than the sharp bound $\underline{{F}}$. This worsening is due to the fact that we need to use Neural Networks that satisfy a slightly worse Lipschitz property, to ensure that we have sufficient approximation power. Although Theorem 4 does not guarantee $\{F_{n}\}$ converges to a point, it states that $\{F_{n}\}$ may oscillate in the interval $[\underline{F}^{\hat{L}_{f}},F_{*}]$, which will still give a useful lower bound to ground-truth value $F_{*}$. In particular, if the graph is identifiable, we have $F^{*}=\underline{F}^{\hat{L}_{f}}$ and $\{F_{n}\}$ converges to $F^{*}$. Also, note that in Theorem 4, we use wide NNs rather than deep NN for $\mathcal{F}_{0,n}$ because results in [52] show that wide NNs can approximate Lipschitz functions while controlling their Lipschitz constants (a property that is not yet established for deep NNs). Similar results can be obtained if wide neural nets are used for all components, invoking Theorem 2.
As a special case, we leverage Theorem 4 for a non-asymptotic rate for the ATE without confounding.
Proposition 3 (Hölder continuity of ATE).
Given two causal models ,^$\mathcal{M},\hat{\mathcal{M}}\in\mathcal{M}(\mathcal{G},\mathcal{F}_{L},\boldsymbol{U})$ satisfying 1 and 3, let their observational distributions be $\nu,\mu$. Suppose the norms of all variables are bounded by $K>0$. If (1) (Overlap) $\nu$ is absolutely continuous with respect to one probability measure $P$ and the density $p_{\nu}\left(t|\text{Pa}(T)=x\right)\geqslant\delta>0$ for x almost surely and $t\in[t_{1},t_{2}]$ and (2) (No Confounding) there is no confounding in the causal graph $\mathcal{G}$, we have
$\int_{t_{1}}^{t_{2}}(\mathbb{E}_{\mathcal{M}}[Y(t)]-\mathbb{E}_{\hat{\mathcal{M}}}[\hat{Y}(t)])^{2}P(dt)$ $\leqslant\frac{2C_{W}}{\delta}W(\mu,\nu)+2(L+1)^{n_{V}}W^{2}(\mu,\nu)(t_{2}-t_{1}),$
@@
Corollary 2.
Let $F,\mu_{T}$ in Theorem 4 to be $F(\boldsymbol{V})=Y,\mu_{T}\sim\text{Unif}([t_{1},t_{2}]),\epsilon>0$. Suppose that the assumptions in Proposition 3 and Theorem 4 hold, with probability at least $1-O(n^{-2})$, we have $|F_{n}-F_{*}|\leqslant O(\sqrt{\alpha_{n}})$, where $F_{n},F_{*},\alpha_{n}$ are defined in Theorem 4.

### v2

@@ 3.2 Approximating Mixed Distributions by Deep Neural Networks
Assumption 3.6 (Lower Bound).
@@
3.6 implies that $d\lambda/d(f^{-1}_{\#}\mathbb{P})$ exists and is lowered bounded by a constant $C_{f}$. The next proposition extends the Skorohod representation theorem [48]. It shows that under 3.6, it is possible to simulate any distribution on the unit cubes with Hölder continuous curves and uniform distribution on $[0,1]$.
Proposition 3.7.
Let $\lambda$ be the Lebesgue measure on $[0,1]$. Given any probability measure $\mathbb{P}$ that satisfies 3.6, there exists a continuous curve $\gamma:[0,1]\rightarrow\text{supp}(\mathbb{P})$ such that $\gamma_{\#}\lambda=\mathbb{P}$. Furthermore, if $d\geqslant 1$, $\gamma$ is $1/d$-Hölder continuous.
Results from [53] show that we can approximate any Hölder continuous $d$-dimensional function with index $\alpha$ by a deep ReLu network with depth $L$ and error $O(L^{-2\alpha/d})$. Leveraging this result, we can replace the first part of the architecture in the previous subsection with deep ReLu networks (See Figure 9 in the appendix). The remaining two parts are the same as the construction in Figure 2. The following theorem gives a sharper bound on the approximation error compared with Theorem 3.5.
Theorem 3.8.
Under the 3.4, if in addition, $\mathbb{P}$ constrained to each component satisfies 3.6, there exists a neural network $\widehat{g}=\widehat{g}_{3}^{\tau}\circ\widehat{g}_{2}\circ{\widehat{g}_{1}}$ with the above architecture such that
$W(\widehat{g}_{\#}\lambda,\mathbb{P})\leqslant O\left(L_{1}^{-2/\max_{i}\{d_{i}^{C}\}}+L_{2}^{-2/\max_{i}\{d_{i}^{C}\}}+(\tau-\tau\log\tau)\right),$
where $d_{i}^{C}$ are the dimensions of the connected components in 3.4, $\widehat{g}_{i},i=1,2$ has the form $\left(\widehat{g}_{i}^{1}(x_{1}),\cdots,\widehat{g}_{i}^{N_{C}}(x_{N_{C}})\right)$ and $\widehat{g}_{1}^{j}\in\mathcal{NN}_{1,d_{j}^{C}}(\Theta(d_{j}^{C}),L_{1})$. $\widehat{g}_{2}^{j},\widehat{g}_{3}^{\tau},\tau$ are the same as in Theorem 3.5.
Let $N$ be the number of nonzero weights in a neural network, Theorem 3.8 shows that a deep NN can achieve $O(N^{-2/d})$ error while by Theorem 3.5 a wide network can only achieve $O(N^{-1/d})$ error. Now, we can put things together to construct NCM approximations. For simplicity, we omit the input and output dimensions of the neural network. As we mention in previous sections, our construction (3.3) consists of two parts, $\widehat{f}_{i}$ approximating the structural functions $f_{i}$ in (3.1), and $\widehat{g}_{j}(Z_{j})=\widehat{g}_{3,j}^{\tau}\circ\widehat{g}_{2,j}\circ\widehat{g}_{1,j}(Z_{j})$ approximating the latent variables $U_{j}$. Let $\text{NCM}_{\mathcal{G}}(\mathcal{F}_{0},\mathcal{F}_{1},\mathcal{F}_{2},\tau)$ be a collection of NCMs with structural equation (3.3) and
$\widehat{f}_{i}\in\mathcal{F}_{0},\widehat{g}_{1,j}=(\widehat{g}^{1}_{1,j},\cdots,\widehat{g}^{N_{C,j}}_{1,j}),\widehat{g}^{i}_{1,j}\in\mathcal{F}_{1},\widehat{g}_{2,j}=(\widehat{g}^{1}_{2,j},\cdots,\widehat{g}^{N_{C,j}}_{2,j}),\widehat{g}^{i}_{2,j}\in\mathcal{F}_{2},$
@@
Corollary 3.9.
Given a causal model $\mathcal{M}^{*}\in\mathcal{M}(\mathcal{G},\mathcal{F}_{L},\bm{U})$, suppose that Assumptions 2.3-3.2 hold and the distributions of $U_{C}$ for all $C^{2}$-component satisfy the assumptions of Theorem 3.8. Let $d_{\max}^{\text{in}}$ and $d_{\max}^{\text{out}}$ be the largest input and output dimension of $f_{i}$ in (3.1) and $d_{\max}^{U}$ be the largest dimension of all latent variables. There exists a neural causal model
^$\widehat{\mathcal{M}}\in\text{NCM}_{\mathcal{G}}(\mathcal{NN}(\Theta(d^{\text{in}}_{\max}d^{\text{out}}_{\max}),L_{0}),\mathcal{NN}\left(\Theta\left(d^{U}_{\max}\right),L_{1}\right),\mathcal{NN}\left(\Theta\left((d^{U}_{\max})^{2}\right),L_{2}\right),\tau)$
with structure equations (3.3). For any intervention $\bm{T}=\bm{t}$, $\widehat{\mathcal{M}}$ satisfies
$W\left(P^{\mathcal{M}^{*}}(\bm{V}(\bm{t})),P^{\widehat{\mathcal{M}}}(\bm{V}(\bm{t}))\right)\leqslant O(L_{0}^{-2/d^{\text{in}}_{\max}}+L_{1}^{-2/d^{U}_{\max}}+L_{2}^{-2/d^{U}_{\max}}+(\tau-\tau\log\tau)).$
@@ 4 Consistency of Neural Causal Partial Identification
In this section, we prove the consistency of the max/min optimization approach to partial identification. In the finite sample setting, we consider the following estimator $F_{n}$ of the optimal values of (1.1).
$F_{n}=\mathop{\mathrm{argmin}}_{\widehat{\mathcal{M}}\in\text{NCM}_{\mathcal{G}}(\mathcal{F}_{0,n},\mathcal{F}_{1,n},\mathcal{F}_{2,n},\tau_{n})}$ $\mathbb{E}_{t\sim\mu_{T}}\mathbb{E}_{\widehat{\mathcal{M}}}[F(V_{1}(t),\cdots,V_{n_{V}}(t))],$ (4.1)
$s.t.~{}~{}~{}$ $S_{\lambda_{n}}(P_{m_{n}}^{\widehat{\mathcal{M}}}(\bm{V}),P^{\mathcal{M}^{*}}_{n}(\bm{V}))\leqslant\alpha_{n},$
where $P^{\mathcal{M}^{*}}_{n},P_{m_{n}}^{\widehat{\mathcal{M}}}$ are the empirical distribution of $P^{\mathcal{M}^{*}},P^{\widehat{\mathcal{M}}}$ with sample size $n,m_{n}$, $\mu_{T}$ is some given measure and $\mathcal{F}_{i,n}$ will be specified later. For example, the counterfactual outcome $\mathbb{E}[Y(1)]$ is a special case of the objective. Our results can be easily generalized to any linear combination of objective functions of this form. We use the Sinkhorn distance because it can be computed efficiently in practice [15]. We want to study if $F_{n}$ gives a useful lower bound as $n\rightarrow\infty$.
@@
Proposition 4.1 (Informal, see Proposition C.1 for a formal version).
There exists a constant $c>0$ and an identifiable SCM $\mathcal{M}^{*}$ satisfying Assumptions 2.3-3.6 such that for any $\epsilon>0$, there exists an SCM $\mathcal{M}_{\epsilon}$ satisfying $W(P^{\mathcal{M}^{*}}(\bm{V}),P^{\mathcal{M}_{\epsilon}}(\bm{V}))\leq\epsilon$ and $|\text{ATE}_{\mathcal{M}^{*}}-\text{ATE}_{\mathcal{M}_{\epsilon}}|>c$.
Here, $\mathcal{M}^{*}$ is the ground-truth model and $\mathcal{M}_{\epsilon}$ are the models we use to approximate $\mathcal{M}^{*}$. Proposition 4.1 implies that we need some regularization on $\mathcal{M}_{\epsilon}$. Otherwise, even if the observation distributions are close, their ATEs can be far away. In particular, we may want to regularize the Lipschitz constant of the NN. Much work has been done to impose Lipschitz regularization during the training process [13, 50, 22, 42, 10]. We denote $\text{Lip}(f)$ to be the Lipschitz constant of a function $f$ and define the truncated Lipschitz NN class,
@@
For simplicity, we omit the dimensions and use shorthand $\mathcal{N}\mathcal{N}^{L_{f},K}(W,L)$. The next theorem gives the consistency result of the min estimator. To state the theorem, we define $F_{*}=\mathbb{E}_{t\sim\mu_{T}}\mathbb{E}_{\mathcal{M}^{*}}[F(V_{1}(t),\cdots,V_{n_{V}}(t))]$ to be the true value and $\underline{F}^{L}$ to be the optimal value of the following optimization problem over SCMs with Lipschitz constant $L$.
$\underline{F}^{L}=\mathop{\mathrm{argmin}}_{\widehat{\mathcal{M}}\in\mathcal{M}\left(\mathcal{G},\mathcal{F}^{K}_{L},\bm{U}\right),P(\bm{U})}$ $\mathbb{E}_{t\sim\mu_{T}}\mathbb{E}_{\widehat{\mathcal{M}}}[F(V_{1}(t),\cdots,V_{n_{V}}(t))],$ (4.2)
$s.t.$ $W(P^{\widehat{\mathcal{M}}}(\bm{V}),P^{\mathcal{M}^{*}}(\bm{V}))=0,$
where the minimum is taken over $\mathcal{M}\left(\mathcal{G},\mathcal{F}^{K}_{L},\bm{U}\right)$ with $\mathcal{F}^{K}_{L}=\{f:\|f\|\leqslant K,f\in\mathcal{F}_{L}\}$ and all latent distributions $P(\bm{U})$. Note that if $L$ is the Lipschitz bound $L_{f}$ that we assume on our structural functions, then $\underline{{F}}^{L}$ is the sharp lower bound.
Theorem 4.2.
Let $\mathcal{M^{*}}$ be any SCM satisfying the assumptions of Corollary 3.9. Suppose that the Lipschitz constant of functions in $\mathcal{M}$ is $L_{f}$, $F:\mathbb{R}^{n_{V}}\rightarrow\mathbb{R}$ in (4.1) is Lipschitz continuous and $\tau_{n}>0$, let $K>0$ be the constant in 3.2, $\widehat{L}_{f}=\sqrt{d_{\max}^{\text{in}}d_{\max}^{\text{out}}}L_{f}$,
$\mathcal{F}_{0,n}=$ $\mathcal{NN}^{\widehat{L}_{f},K}\left(W_{0,n},\Theta\left(\log d^{\text{in}}_{\max}\right)\right),\mathcal{F}_{1,n}=\mathcal{NN}^{\infty,\infty}(\Theta(d^{U}_{\max}),L_{i,n}),$
@@
where $d_{\max}^{U},d_{\max}^{\text{in}}$ and $d_{\max}^{\text{out}}$ are defined in Corollary 3.9, take the radius to be $\alpha_{n}=\epsilon_{n}+s_{n},\epsilon_{n}=O(W_{0,n}^{-1/d^{\text{in}}_{\max}}+\sum_{i=1}^{2}L_{i,n}^{-2/d^{U}_{\max}}+\tau_{n}\log\tau_{n}),$
$s_{n}=O(m_{n}^{-1/\left(d_{\max}^{U}+2\right)}\log m_{n}+\delta_{n}+\log(nm_{n})\lambda_{n}),\quad\delta_{n}=O(n^{-1/{\max\{2,d_{\max}^{U}\}}}\log^{2}(n)).$
If $m_{n}=\Omega(n),L_{i,n}=\Theta(m_{n}^{d_{\max}^{U}/\left(2d_{\max}^{U}+4\right)}),i=1,2,\lim_{n\rightarrow\infty}\min\{\tau_{n}^{-1},W_{0,n},(\log(nm_{n})\lambda_{n})^{-1}\}=\infty$ , then with probability 1, $[\liminf_{n\rightarrow\infty}F_{n},\limsup_{n\rightarrow\infty}F_{n}]\subset[\underline{F}^{\widehat{L}_{f}},F_{*}]$.
Note that our theorem shows that the lower limit of the solution is large than the point $\underline{{F}}^{\widehat{L}_{f}}$, where $\widehat{L}_{f}$ is slightly larger than the original constraint $L_{f}$ we impose on the structural functions. Hence, this point can potentially be slightly smaller than the sharp bound $\underline{{F}}^{L_{f}}$. This worsening is due to the fact that we need to use NNs that satisfy a slightly worse Lipschitz property, to ensure that we have sufficient approximation power. Although Theorem 4.2 does not guarantee $\{F_{n}\}$ converges to a point, it states that $\{F_{n}\}$ may oscillate in the interval $[\underline{F}^{\widehat{L}_{f}},F_{*}]$, which will still give a useful lower bound to ground-truth value $F_{*}$. In particular, if the graph is identifiable, we have $F^{*}=\underline{F}^{\widehat{L}_{f}}$ and $\{F_{n}\}$ converges to $F^{*}$. Also, note that in Theorem 4.2, we use wide NNs rather than deep NN for $\mathcal{F}_{0,n}$ because results in [53] show that wide NNs can approximate Lipschitz functions while controlling their Lipschitz constants (a property that is not yet established for deep NNs). Similar results can be obtained if wide neural nets are used for all components, invoking Theorem 3.5.
As a special case, we leverage Theorem 4.2 for a non-asymptotic rate for the ATE without confounding.
Proposition 4.3 (Hölder continuity of ATE).
Given two causal models ,^$\mathcal{M},\widehat{\mathcal{M}}\in\mathcal{M}(\mathcal{G},\mathcal{F}_{L},\bm{U})$ satisfying 2.3 and 3.2, let their observational distributions be $\nu,\mu$. Suppose the norms of all variables are bounded by $K>0$. If (1) (Overlap) $\nu$ is absolutely continuous with respect to one probability measure $P$ and the density $p_{\nu}\left(t|\text{Pa}(T)=x\right)\geqslant\delta>0$ for x almost surely and $t\in[t_{1},t_{2}]$ and (2) (No Confounding) there is no confounding in the causal graph $\mathcal{G}$, we have
$\int_{t_{1}}^{t_{2}}(\mathbb{E}_{\mathcal{M}}[Y(t)]-\mathbb{E}_{\widehat{\mathcal{M}}}[\widehat{Y}(t)])^{2}P(dt)$ $\leqslant\frac{2C_{W}}{\delta}W(\mu,\nu)+2(L+1)^{n_{V}}W^{2}(\mu,\nu)(t_{2}-t_{1}),$
@@
Corollary 4.4.
Let $F,\mu_{T}$ in Theorem 4.2 to be $F(\bm{V})=Y,\mu_{T}\sim\text{Unif}([t_{1},t_{2}]),\epsilon>0$. Suppose that the assumptions in Proposition 4.3 and Theorem 4.2 hold, with probability at least $1-O(n^{-2})$, we have $|F_{n}-F_{*}|\leqslant O(\sqrt{\alpha_{n}})$, where $F_{n},F_{*},\alpha_{n}$ are defined in Theorem 4.2.